# ReGlove: A Soft Pneumatic Glove for Activities of Daily Living Assistance via Wrist-Mounted Vision

### v1

@@ Abstract
This paper presents ReGlove, a system that converts low-cost commercial pneumatic rehabilitation gloves into vision-guided assistive orthoses. Chronic upper-limb impairment affects millions worldwide, yet existing assistive technologies remain prohibitively expensive or rely on unreliable biological signals. Our platform integrates a wrist-mounted camera with an edge-computing inference engine (Raspberry Pi 5) to enable context-aware grasping without requiring reliable muscle signals. By adapting real-time YOLO-based computer vision models, the system achieves grasp classification accuracy with sub- end-to-end latency. Physical validation using standardized benchmarks shows success on YCB object manipulation and reliable performance across Activities of Daily Living (ADL) tasks. With a total cost under $ and exclusively commercial components, ReGlove provides a technical foundation for accessible, vision-based upper-limb assistance that could benefit populations excluded from traditional EMG-controlled devices.
@@ I INTRODUCTION
Recent advances in computer vision for prosthetic control demonstrate that visual context can robustly inform grasp selection [degol2016, taverne2019]. However, these approaches have not been systematically applied to orthotic applications using commercial components. The ReGlove system bridges this gap by integrating established computer vision techniques with affordable, commercially available hardware.
This paper presents three key contributions: (1) An integrated hardware-software architecture that transforms commercial pneumatic gloves into vision-guided orthoses using readily available components; (2) A lightweight perception pipeline based on YOLO architectures that achieves real-time grasp classification on edge computing hardware; and (3) A comprehensive performance evaluation establishing baseline functionality across standardized benchmarks including YCB object manipulation and ADL tasks. Through this proof-of-concept, we demonstrate a viable pathway toward assistive devices that balance capability with accessibility.
@@ II-A Actuation for Hand Assistance
Hand assistive devices primarily employ cable-driven or pneumatic actuation. Cable-driven systems [kim2025, rose2019] transmit force from proximal motors through tendon-like mechanisms, offering precise control but suffering from mechanical complexity, cable management issues, and limited compliance.
Pneumatic actuators, used in commercial rehabilitation gloves, provide inherent compliance and safety through soft, inflatable chambers [lim2023]. Clinical evidence supports their efficacy in improving hand function, with randomized trials showing significant improvements in active range of motion and grip strength for chronic stroke patients [fardipour2022, ko2023]. Their commercial availability and low cost (<$) make them a practical foundation for accessible assistive technology.
Alternative approaches include shape-memory alloys [butzer2021] and motorized exoskeletons, but these face challenges in reliability, weight, and cost that limit practical deployment.
@@ II-B Control Modalities
Traditional control methods include manual triggers and surface electromyography (sEMG). Manual control requires use of the contralateral limb, making it impractical for independent use. sEMG-based control can enable more natural actuation but often fails for patients with weak or noisy signals due to neuromuscular degeneration [xu2025].
Vision-based control, successfully demonstrated in prosthetic systems [degol2016, zhang2025], offers a promising alternative by relying on object context rather than biological signals. Prior work primarily used computationally intensive architectures like VGG-16, limiting real-time performance on low-power hardware. We adapt this approach using modern YOLO architectures optimized for edge deployment, making vision-based control practical for orthotic applications where EMG may be unreliable.
@@ III-A Hardware Implementation
The pneumatic subsystem uses a commercial rehabilitation glove with ethylene-vinyl acetate (EVA) bellows actuators, providing one degree of freedom per finger for bidirectional flexion and extension. We employ two HG095 mini air pumps ( flow rate) for inflation and vacuum generation, and six ZHV-0519 three-way solenoid valves for individual finger control.
@@
Thumb Adaptation: The commercial glove’s single-DOF design limits thumb opposition. We address this with a custom 3D-printed thermoplastic polyurethane (TPU) brace that maintains partial abduction while allowing pneumatic flexion, preserving capability for most functional grasp types [zheng2011].
@@
Control Inputs: While the system architecture supports multiple input modalities (sEMG, EEG, EOG), we use a simple tactile switch for benchtop validation to isolate vision system performance. This allows future drop-in replacement with sEMG once IRB approval is secured for clinical studies.
@@
Note: Costs are approximate and vary based on supplier.
@@ III-B Vision Pipeline & Model Development
We used a grasp classification system using three publicly available datasets: DeepGrasping ( images) [degol2016], ImageNet subset ( images), and HandCam ( images) [taverne2019]. To address class imbalance, we applied extensive data augmentation including geometric transformations, photometric adjustments, and occlusion modeling, yielding approximately images per grasp type (pinch, power, three-jaw chuck, tool, key).
@@
VGG-16 & VGG-16 + Depth: Baseline models replicating prior work [degol2016]
@@
Depth augmentation using synthetic depth maps from DepthAnything [yang2024] did not improve performance, likely due to inconsistency in synthetic depth quality. Both YOLO variants significantly outperformed VGG-based approaches (Table II), with YOLO v11 achieving accuracy versus for VGG-16. YOLO’s superior performance stems from architectural features that preserve spatial structure (SPPF, FPN/PAN layers) and integrated augmentation mechanisms that improve robustness to lighting and background variation.
Given its optimal accuracy-latency tradeoff, we selected YOLO v11 for system integration, achieving inference latency on Raspberry Pi 5—well below the threshold for human-perceptible feedback [jerald2009relatingscene].
@@ IV-B1 YCB Object Set
Using the YCB Gripper Assessment Protocol [calli2015], ReGlove achieved an overall success rate of (/ points). Performance was robust for objects with defined edges and surfaces (cups, blocks, utensils) but lower for small, smooth, or low-friction items (marbles, coins, washers). This performance gap primarily reflects mechanical limitations of the compliant EVA actuators rather than perception errors. Full results are available in supplementary materials (Table S-III).
@@ IV-B2 Activities of Daily Living (ADL)
On a subset of ADL tasks based on Matheus & Dollar [matheus2010], the system achieved a mean performance score of out of (=failed, =excellent). The system excelled at tasks involving power or tripod grasps (pouring liquids, manipulating utensils) but struggled with fine manipulation requiring precise fingertip control (unwrapping tablets, rotating small bolts).
@@ V-A Technical Performance and Significance
The system’s grasp classification accuracy and end-to-end latency compare favorably with prior vision-based prosthetic systems requiring more complex hardware [degol2016, taverne2019]. More significantly, by relying exclusively on visual context rather than biological signals, the approach extends accessibility to patient populations with unreliable EMG due to neuromuscular degeneration [xu2025].
@@ V-C Future Directions
Multi-modal control integration: Subsequent iterations will incorporate surface electromyography (sEMG) as the primary intent detection modality, operating in concert with the existing vision-based grasp classification. This hybrid approach will enable more natural actuation paradigms while maintaining the robustness of visual context awareness. Additionally, implementation of closed-loop force control will enhance manipulation precision and user experience.
@@
Clinical translation: Formal studies with stroke and SCI patients to quantify ADL improvement, user acceptance, and long-term usability.

### v2

@@ Abstract
This paper presents ReGlove, a system that converts low-cost commercial pneumatic rehabilitation gloves into vision-guided assistive orthoses. Chronic upper-limb impairment affects millions worldwide, yet existing assistive technologies remain prohibitively expensive or rely on unreliable biological signals. Our platform integrates a wrist-mounted camera with an edge-computing inference engine (Raspberry Pi 5) to enable context-aware grasping without requiring reliable muscle signals. By adapting real-time You-Only-Look-Once (YOLO) based computer vision models, the system achieves grasp classification accuracy with sub- end-to-end latency. Physical validation using standardized benchmarks shows success on Yale-CMU-Berkeley (YCB) Object and Model Set object manipulation and reliable performance across Activities of Daily Living (ADL) tasks. With a total cost under $ and exclusively commercial components, ReGlove provides a technical foundation for accessible, vision-based upper-limb assistance that could benefit populations excluded from traditional Electromyography (EMG)-controlled devices.
@@ I INTRODUCTION
Recent advances in computer vision for prosthetic control demonstrate that visual context can robustly inform grasp selection [3, 11]. However, these approaches have not been systematically applied to orthotic applications using commercial components. The ReGlove system bridges this gap by integrating established computer vision techniques with affordable, commercially available hardware.
This paper presents three key contributions: (1) An integrated hardware-software architecture that transforms commercial pneumatic gloves into vision-guided orthoses using readily available components; (2) A lightweight perception pipeline based on YOLO architectures that achieves real-time grasp classification on edge computing hardware; and (3) A comprehensive performance evaluation establishing baseline functionality across standardized benchmarks including YCB object manipulation and Activities of Daily Living (ADL) tasks. Through this proof-of-concept, we demonstrate a viable pathway toward assistive devices that balance capability with accessibility.
@@ II-A Actuation for Hand Assistance
Hand assistive devices primarily employ cable-driven or pneumatic actuation. Cable-driven systems [6, 10] transmit force from proximal motors through tendon-like mechanisms, offering precise control but suffering from mechanical complexity, cable management issues, and limited compliance.
Pneumatic actuators, used in commercial rehabilitation gloves, provide inherent compliance and safety through soft, inflatable chambers [8]. Clinical evidence supports their efficacy in improving hand function, with randomized trials showing significant improvements in active range of motion and grip strength for chronic stroke patients [4, 7]. Their commercial availability and low cost (<$) make them a practical foundation for accessible assistive technology.
Alternative approaches include shape-memory alloys [1] and motorized exoskeletons, but these face challenges in reliability, weight, and cost that limit practical deployment.
@@ II-B Control Modalities
Traditional control methods include manual triggers and sEMG. Manual control requires use of the contralateral limb, making it impractical for independent use. sEMG-based control can enable more natural actuation but often fails for patients with weak or noisy signals due to neuromuscular degeneration [12].
Vision-based control, successfully demonstrated in prosthetic systems [3, 14], offers a promising alternative by relying on object context rather than biological signals. Prior work primarily used computationally intensive architectures like VGG-16, limiting real-time performance on low-power hardware. We adapt this approach using modern YOLO architectures optimized for edge deployment, making vision-based control practical for orthotic applications where EMG may be unreliable.
@@ III-A Hardware Implementation
The pneumatic subsystem uses a commercial rehabilitation glove with ethylene-vinyl acetate bellows actuators, providing one degree of freedom per finger for bidirectional flexion and extension. We employ two HG095 mini air pumps ( flow rate) for inflation and vacuum generation, and six ZHV-0519 three-way solenoid valves for individual finger control. The complete wiring schematic is shown in Fig. 1.
@@
Thumb Adaptation: The commercial glove’s single degrees of freedom design limits thumb opposition. We address this with a custom 3D-printed thermoplastic polyurethane brace that maintains partial abduction while allowing pneumatic flexion, preserving capability for most functional grasp types [15]. The hand configuration with and without the brace and glove is shown in Fig. 4.
@@
Control Inputs: While the system architecture supports multiple input modalities (surface Electromyography, Electroencephalography, Electrooculography), we use a simple tactile switch for benchtop validation to isolate vision system performance. This allows future drop-in replacement with sEMG once IRB approval is secured for clinical studies.
@@ III-B Vision Pipeline & Model Development
We used a grasp classification system using three publicly available datasets: DeepGrasping ( images) [3], ImageNet subset ( images), and HandCam ( images) [11]. To address class imbalance, we applied extensive data augmentation including geometric transformations, photometric adjustments, and occlusion modeling, yielding approximately images per grasp type (pinch, power, three-jaw chuck, tool, key).
@@
VGG-16 & VGG-16 + Depth: Baseline models replicating prior work [3]
@@
Depth augmentation using synthetic depth maps from DepthAnything [13] did not improve performance, likely due to inconsistency in synthetic depth quality. Both YOLO variants significantly outperformed VGG-based approaches (Table II), with YOLO v11 achieving accuracy versus for VGG-16. YOLO’s superior performance stems from architectural features that preserve spatial structure (Spatial Pyramid Pooling - Fast, Feature Pyramid Network/Path Aggregation Network layers) and integrated augmentation mechanisms that improve robustness to lighting and background variation.
Given its optimal accuracy-latency tradeoff, we selected YOLO v11 for system integration, achieving inference latency on Raspberry Pi 5—well below the threshold for human-perceptible feedback [5].
@@ IV-B1 YCB Object Set
Using the YCB Gripper Assessment Protocol [2], ReGlove achieved an overall success rate of (/ points). Performance was robust for objects with defined edges and surfaces (cups, blocks, utensils) but lower for small, smooth, or low-friction items (marbles, coins, washers). This performance gap primarily reflects mechanical limitations of the compliant ethylene-vinyl acetate actuators rather than perception errors. Full results are available in supplementary materials (Table S-III).
@@ IV-B2 Activities of Daily Living (ADL)
On a subset of daily living tasks based on Matheus & Dollar [9], the system achieved a mean performance score of out of (=failed, =excellent). The system excelled at tasks involving power or tripod grasps (pouring liquids, manipulating utensils) but struggled with fine manipulation requiring precise fingertip control (unwrapping tablets, rotating small bolts). A comparative analysis of execution times for these tasks is shown in Fig. 3.
@@ V-A Technical Performance and Significance
The system’s grasp classification accuracy and end-to-end latency compare favorably with prior vision-based prosthetic systems requiring more complex hardware [3, 11]. More significantly, by relying exclusively on visual context rather than biological signals, the approach extends accessibility to patient populations with unreliable EMG due to neuromuscular degeneration [12].
@@ V-C Future Directions
Multi-modal control integration: Subsequent iterations will incorporate sEMG as the primary intent detection modality, operating in concert with the existing vision-based grasp classification. This hybrid approach will enable more natural actuation paradigms while maintaining the robustness of visual context awareness. Additionally, implementation of closed-loop force control will enhance manipulation precision and user experience.
@@
Clinical translation: Formal studies with stroke and SCI patients to quantify daily living tasks improvement, user acceptance, and long-term usability.
@@ VI CONCLUSION
FUNDING
This research received no external grants or funding. It was supported by the authors’ personal resources.